# for Scalable and Intelligent Embodied Systems

### v2

@@ Abstract
We explore how scalable robot data can address real-world challenges for generalized robotic manipulation. Introducing AgiBot World, a large-scale platform comprising over 1 million trajectories across 217 tasks in five deployment scenarios, we achieve an order-of-magnitude increase in data scale compared to existing datasets. Accelerated by a standardized collection pipeline with human-in-the-loop verification, AgiBot World guarantees high-quality and diverse data distribution. It is extensible from grippers to dexterous hands and visuo-tactile sensors for fine-grained skill acquisition. Building on top of data, we introduce Genie Operator-1 (GO-1), a novel generalist policy that leverages latent action representations to maximize data utilization, demonstrating predictable performance scaling with increased data volume. Policies pre-trained on our dataset achieve an average performance improvement of 30% over those trained on Open X-Embodiment, both in in-domain and out-of-distribution scenarios. GO-1 exhibits exceptional capability in real-world dexterous and long-horizon tasks, achieving over 60% success rate on complex tasks and outperforming prior RDT approach by 32%. By open-sourcing the dataset, tools, and models, we aim to democratize access to large-scale, high-quality robot data, advancing the pursuit of scalable and general-purpose intelligence. This work is a collaborative effort between Shanghai AI Lab, AgiBot Inc. and Shanghai Innovation Institute. For detailed authorship, please visit the official website and Github repo.
@@ I Introduction
Dataset Traj. Skill Scene Detailed Annotation Cam. Calibration Arm Type Dex. Hand Failure Recovery Human-in- the-loop Collection
@@
Contribution. 1) We construct AgiBot World dataset, a multifarious robot learning dataset accompanied by open-source tools to advance research on policy learning at scale. As a pioneering initiative, AgiBot World employs an inclusive optimized pipeline, from scene configuration, task design, data collection, to human-in-the-loop verification, which ensures unparalleled data quality. 2) We propose GO-1, a robot foundation policy using latent action representations to unlock web-scale pre-training on heterogeneous data. Empowered by AgiBot World dataset, it outperforms prior generalist policies in generalization and dexterity, with performance scaling predictably with dataset size.
@@ II Related Work
Data scaling in robotics. Robot learning datasets from automated scripts or human teleoperation have enabled policy learning, with early efforts like RoboTurk [19] and BridgeData [12] offering small-scale datasets with 2.1k and 7.2k trajectories, respectively. Larger datasets, such as RT-1 [14] (130k trajectories), expand scopes yet remain limited to few environments and skills. Open X-Embodiment [6] aggregates various datasets into a unified format, growing to more than 2.4 million trajectories, as a consequence it suffers from significant variability in embodiments, observation perspectives, and inconsistent data quality, limiting its overall effectiveness. More recently, DROID [7] moves towards scaling up scenes for greater diversity by crowd-sourcing demonstrations yet falls short in data scale and quality control. Prior datasets above generally face limitations in data scale, task practicality, and scenario naturalness, compounded by inadequate quality assurance and hardware restrictions, which impedes generalist policy training. As shown in Tab. I, our dataset addresses these gap adequately. We build a data collection facility spanning five scenarios to reconstruct real-world diversity and authenticity. With over 1 million trajectories gathered by skilled teleoperators through rigorous verification protocols, AgiBot World utilizes humanoid robots equipped with visuo-tactile sensors and dexterous hands to enable multimodal demonstrations, setting it apart from previous efforts. Unlike Pumacay et al. [20], which serves as a simulation benchmark for evaluating generalization, what we propose is a full-stack platform with data, models, benchmarks, and ecosystem.
@@ IV-A Latent Action Model
To extract latent actions from video frames $\{I_{t},I_{t+H}\}$, the latent action model is constructed around an inverse dynamics model-based encoder $\mathbf{I}(z_{t}|I_{t},I_{t+H})$ and a forward dynamics model-based decoder $\mathbf{F}(I_{t+H}|I_{t},z_{t})$. The encoder employs a spatial-temporal transformer [31] with casual temporal masks following Bruce et al. [30], while the decoder is a spatial transformer that takes the initial frame and discretized latent action tokens $z_{t}=[z_{t}^{0},...,z_{t}^{k-1}]$ as input, with $k$ set to 4. The latent action tokens are quantized using a VQ-VAE objective [32], with a codebook of size $|C|$.
@@ V-A2 Implementation Details
The AgiBot World alpha represents the partial subset of our dataset, constituting approximately 14% of the full-version, AgiBot World beta (a.k.a. last row in Tab. I). Following the completion of the third-stage pre-training, the pre-trained GO-1 exhibits basic competency in task completion. Unless otherwise specified, we further enhance the model by fine-tuning it using high-quality, task-specific demonstrations, enabling adaptation to new tasks for evaluation. For GO-1, fine-tuning is conducted with a learning rate of 2e-5, a batch size of 768, and 30,000 optimization steps.
@@ V-C Is GO-1 a more capable generalist policy?
We evaluate GO-1 on five tasks of varying complexity, categorized by their visual richness and task horizon. The results, as shown in Fig. 5, are averaged over 30 trials per task, with 10 trials conducted in a seen setup and 20 trials under variations or distractions. GO-1 significantly outperforms RDT, particularly in tasks such as “Pour Water”, which demands robustness to object positions, and “Restock Beverage”, which requires visual robustness and instruction-following capabilities. The inclusion of the latent planner in the ViLLA model further improves performance, resulting in an average improvement of 0.12 task completion score.
@@ V-E How does data quality impact policy learning?
We explore the impact of quality checks introduced in our human-in-the-loop data collection on policy learning. Specifically, we provide an ablation study by fine-tuning an RDT model pre-trained on the alpha dataset using both verified (528 trajectories) and unverified (482 trajectories) data from the “Wipe Table” task. As shown in Fig. 7(b), being larger in quantity does not necessarily translate to improved performance, while a smaller set of human-verified data yields a 0.18 boost in the completion score, underscoring the importance of high-quality data for policy learning.
@@ Acknowledgement
Qingwen Bu, Guanghui Ren, Chiming Liu, Chengen Xie,
Modi Shi, Xindong He, Jianheng Song,
@@
Jianlan Luo, Jiangmiao Pang, Bin Zhao,
Junchi Yan, Ping Luo
@@
How to cite
  • Github repository
In the technical report, we list the authors in alphabetical order by their last names.

### v4

@@ Abstract
We explore how scalable robot data can address real-world challenges for generalized robotic manipulation. Introducing AgiBot World, a large-scale platform comprising over 1 million trajectories across 217 tasks in five deployment scenarios, we achieve an order-of-magnitude increase in data scale compared to existing datasets. Accelerated by a standardized collection pipeline with human-in-the-loop verification, AgiBot World guarantees high-quality and diverse data distribution. It is extensible from grippers to dexterous hands and visuo-tactile sensors for fine-grained skill acquisition. Building on top of data, we introduce Genie Operator-1 (GO-1), a novel generalist policy that leverages latent action representations to maximize data utilization, demonstrating predictable performance scaling with increased data volume. Policies pre-trained on our dataset achieve an average performance improvement of 30% over those trained on Open X-Embodiment, both in in-domain and out-of-distribution scenarios. GO-1 exhibits exceptional capability in real-world dexterous and long-horizon tasks, achieving over 60% success rate on complex tasks and outperforming prior RDT approach by 32%. By open-sourcing the dataset, tools, and models, we aim to democratize access to large-scale, high-quality robot data, advancing the pursuit of scalable and general-purpose intelligence. This work is a joint effort among The University of Hong Kong, AgiBot Inc., Shanghai Innovation Institute and Shanghai AI Lab. For detailed authorship, please consult our website and GitHub repository.
@@ I Introduction
Dataset Traj. Skill Scene
Detailed
Annotation
Cam.
Calibration
Arm Type
Dex.
Hand
Failure
Recovery
Human-in-
the-loop
Collection
@@
Contribution. 1) We construct AgiBot World dataset, a multifarious robot learning dataset accompanied by open-source tools to advance research on policy learning at scale. As a pioneering initiative, AgiBot World employs an inclusive optimized pipeline, from scene configuration, task design, data collection, to human-in-the-loop verification, which ensures unparalleled data quality. 2) We propose GO-1, a robot foundation policy using latent action representations to unlock web-scale pre-training on web data. Empowered by AgiBot World dataset, it outperforms prior generalist policies in generalization and dexterity.
@@ II Related Work
Data scaling in robotics. Robot learning datasets from automated scripts or human teleoperation have enabled policy learning, with early efforts like RoboTurk [19] and BridgeData [12] offering small-scale datasets with 2.1k and 7.2k trajectories, respectively. Larger datasets, such as RT-1 [14] (130k trajectories), expand scopes yet remain limited to few environments and skills. Open X-Embodiment [6] aggregates various datasets into a unified format, growing to more than 1.4 million trajectories, as a consequence it suffers from significant variability in embodiments, observation perspectives, and inconsistent data quality, limiting its overall effectiveness. More recently, DROID [7] moves towards scaling up scenes for greater diversity by crowd-sourcing demonstrations yet falls short in data scale and quality control. Prior datasets above generally face limitations in data scale, task practicality, and scenario naturalness, compounded by inadequate quality assurance and hardware restrictions, which impedes generalist policy training. As shown in Tab. I, our dataset addresses these gap adequately. We build a data collection facility spanning five scenarios to reconstruct real-world diversity and authenticity. With over 1 million trajectories gathered by skilled teleoperators through rigorous verification protocols, AgiBot World utilizes humanoid robots equipped with visuo-tactile sensors and dexterous hands to enable multimodal demonstrations, setting it apart from previous efforts. Unlike Pumacay et al. [20], which serves as a simulation benchmark for evaluating generalization, what we propose is a full-stack platform with data, models, benchmarks, and ecosystem.
@@ IV-A Latent Action Model
[…]
@@ V-A2 Implementation Details
The AgiBot World alpha dataset is an early-stage subset, containing partial tasks and roughly 14% of the trajectories in the full beta version. (a.k.a. last row in Tab. I). Following the completion of the third-stage pre-training, the pre-trained GO-1 exhibits basic competency in task completion. Unless otherwise specified, we further enhance the model by fine-tuning it using high-quality, task-specific demonstrations, enabling adaptation to new tasks for evaluation. For GO-1, fine-tuning is conducted with a learning rate of 2e-5, a batch size of 768, and 30,000 optimization steps.
@@ V-C Is GO-1 a more capable generalist policy?
We evaluate GO-1 on five tasks of varying complexity, categorized by their visual richness and task horizon. The results, as shown in Fig. 5, are averaged over 30 trials per task, with 10 trials conducted in a seen setup and 20 trials under variations or distractions. GO-1 significantly outperforms RDT and $\pi_{0}$, particularly in tasks such as “Pour Water”, which demands robustness to object positions, and “Restock Beverage”, which highlights instruction-following capabilities. The inclusion of the latent planner yields an average improvement of 0.12 task completion score.
@@ V-E How does data quality impact policy learning?
We explore the impact of quality checks introduced in our human-in-the-loop data collection on policy learning. Specifically, we provide an ablation study by fine-tuning an RDT model using both verified (528 trajectories) and unverified (482 trajectories) data from the “Wipe Table” task. Verification refers to our “human-in-the-loop” quality assurance method. As shown in Fig. 7(b), being larger in quantity does not necessarily translate to improved performance, while a smaller set of human-verified data yields a 0.18 boost in the completion score, underscoring the importance of high-quality data for policy learning.
@@ Acknowledgement
Qingwen Bu, Guanghui Ren, Chiming Liu, Modi Shi,
Chengen Xie, Xindong He, Jianheng Song,
@@
Jianlan Luo, Bin Zhao, Junchi Yan, Ping Luo
@@
July 2025: IROS camera-ready update
  • We conduct additional comparisons with $\pi_{0}$.